# FingerTip 20K: A Benchmark for Proactive and Personalized Mobile LLM Agents

### v1

@@ 4.3 Data statistics
The summary of data statistics is presented in Table 1. Additionally, Figure 4 reports the distribution of user intent length, episode length, intent categories, and app name in all data. The intent category of each intent is determined by DeepSeek-V3 [22].
@@ 4.4 Personalized action analysis
Table 2: Actions and Format.
Action Format
@@
content=”
type text=”
@@
direction=’down or up
or right or left’
navigate_back ()
navigate_home ()
navigate_recent ()
wait ()
finish ()
Figure 5: Personalized action analysis.
To verify the personalized differences in actions among different types of users, we first simply classified users into different categories based on age groups. Then, we randomly sampled one piece of data from each of the 40 intent categories. For the action sequence of such a piece of data, we calculated the Levenshtein similarity with the action sequence of the most intent-similar data from (i) the same user, (ii) the same type of users, and (iii) different type of users. All similarities were normalized to [0, 100] and plotted in Figure 5. It can be seen that even when performing similar intents, the similarity of action sequences with different types of users is significantly lower than that of the same user or the same type of users.
4.5 Data splits
Table 3: Details on FingerTip train, validation and test splits.
Split # Episodes # Screens # Apps # Categories
Train 16000 177674 460 40
Vali 4411 32859 41 0
Test-suggestion 1000 10412 155 38
Test-execution 200 2074 68 31
We created a training set, a validation set, and two test sets. The number of episodes and features in these sets are detailed in Table 3. Please note that the two test sets contain partially overlapping episodes. The test sets were formed by randomly sampling the last 20% of the data of each user, sorted by time, and then concatenated to ensure coverage of all users and that the proportion of data from each user in the test sets is equal to their proportion in all data. These test sets were used in subsequent experiments.
@@ 5 Experiments
We conducted experiments on some generalist models and some GUI-control agents built on specifically designed models, evaluating their capabilities on the two tracks proposed in the FingerTip benchmark and assessing their performance under different task difficulties. Additionally, we fine-tuned a model using the collected data.
@@ 5.1 Experimental setup
The LLMs we experiment with in this track include GPT-4.1, Qwen-VL-Max, DeepSeek-VL2 [23] and Qwen-2.5-VL-7B [24]. We set the temperature to zero for all models. For proactive task suggestion, the agent only needs one query to output the predicted intent. Since this is a brand new track proposed in our benchmark, there is no mature agent design available for direct use. We have designed a simple prompt to provide to all models evaluated in this track. This prompt contains all necessary inputs (see Section 3.1).
@@
In this track, in addition to the four generalist models mentioned above, we also experiment with three GUI-control agents built on specifically designed models, including Aguvis-7B [16], CogAgent-9B [15] and UI-TARS-1.5-7B [14]. For personalized task execution, the agent needs to interact with the environment in multiple steps to fulfill the user’s instructions. We connect a physical phone to the computer via USB and use Android Debug Bridge (ADB) to provide this environment. Using an emulator would be a more convenient approach, but due to strict app control measures, most Chinese apps can only run on physical phones rather than emulators. For the generalist models, we designed a simple prompt to guide their output of the next action, with the action space consistent with Table 2. This prompt contains all necessary inputs (see Section 3.2). For the three GUI-control agents, they have specific format requirements for input and output. To ensure normal output effects, their original prompts were used, and the input information in Section 3.2 was uniformly integrated into these original prompts. Their output was converted into a form consistent with the action space.
@@
In proactive task suggestion track, the goal of the agent is to maximize the textual similarity between the output and the user’s true intent. We use a pre-trained model, paraphrase-multilingual-MiniLM-L12-v2 [25], to convert the agent’s output and the user’s true intent into embedding vectors and calculate their cosine similarity $S_{1}$. And, we calculate the Levenshtein similarity $S_{2}$ of these two strings. Both similarities are normalized to the range of [0, 1]. Finally, we take $Sim_{1}=(S_{1}+S_{2})/2$ to comprehensively represent the text similarity. In addition to this numerical metric, we also use DeepSeek-V3 [22] to directly determine whether the agent’s output and the user’s true intent can be regarded as the same intent and provide a binary value to evaluate whether the agent successfully predicted the user’s intent, thereby calculating the success rate $SR_{1}$.
In personalized task execution track, the primary goal of the agent is to successfully execute the user’s instructions. We calculate the final success rate $SR_{2}$ by manually checking whether the environment state when the agent outputs $finished()$ matches the user’s instructions. In addition, when the agent steps exceed 2.5 times the golden steps, the task is automatically considered a failure. Note that the path to successfully execute the user’s instructions is not unique. The agent should also make the action sequence reflect the user’s action preferences as much as possible. We do not require the agent’s action at each step to be exactly the same as the user’s golden action. Instead, we calculate the Levenshtein similarity $S_{I}$ of the agent’s complete action sequence and the user’s complete action sequence as two strings. Then, following the approach in Section 4.4, we take the data from the different type of users that is most similar to the current user’s intent, and calculate the Levenshtein similarity $S_{\text{II}}$ of the agent’s complete action sequence and this data’s complete action sequence. Finally, we take the value $Sim_{2}=S_{I}/S_{\text{II}}$. It is obvious that the larger this value is, the more similar the agent’s action sequence is to that of the current user, and the more different it is from that of different type of users. In addition, we measure execution efficiency by comparing the agent steps with the user’s golden steps to calculate the step ratio when the agent successfully execute the user’s instructions. For the two tracks, we also tallied the average time and token count consumed per query to assess the model’s cost.
@@ 5.2 Overall Performance
The overall performance of the models we evaluated in proactive task suggestion is shown in Table 4. Note that here we set the number of $O$ (the first few screenshots of the user’s current actions) to 0. This makes the task quite challenging. GPT-4.1 performed the best among the generalist models, achieving $SR_{1}=7.2$ and $Sim_{1}=0.35$, but it also took the longest time. From $SR_{1}$, it can be intuitively seen that the success rate of all models in predicting the user’s intent is very low, less than 10%. We consider it normal. The models almost only rely on 20 previous intents of the user to predict the user’s current intent, which is difficult to achieve without prior knowledge of the user’s characteristics and intent patterns.
Table 4: Overall performance of proactive task suggestion. Bold indicates the best performance.
@@
Table 5: Overall performance of personalized task execution. Bold indicate the best performance.
@@
The overall performance of the models we evaluated in personalized task execution is shown in Table 5. GPT-4.1 and UI-TARS-1.5-7B achieved the best performance among the generalist models and GUI-control models respectively. The $SR_{2}$ of the generalist models were all very low, mainly due to their lack of precise GUI grounding ability, which led to incorrect UI coordinates being output even when they could correctly analyze the next action, thus failing to interact with the environment accurately. In contrast, the GUI-control models, having undergone targeted training, had stronger abilities to execute instructions and interact with the UI environment, resulting in higher $SR_{2}$, with UI-TARS-1.5-7B having the highest at 38.5. However, the $Sim_{2}$ of all models were approximately 1, indicating that the agent’s action sequence did not favor either the current user or different types of users. This might suggest that the agent tends to complete tasks in a general way without catering to the specific action preferences of users, thus failing to complete tasks in a personalized manner.
@@ 5.3 Effect of task difficulty
Figure 6: Performance under different task difficulties.
@@
For personalized task execution (see Figure 6.b), we calculated the $SR_{2}$ of GUI-control models on different subsets of action lengths in the test set (we did not calculate the $SR_{2}$ of generalist models because their $SR_{2}$ were all too low). It can be seen that as the action length increases, the $SR_{2}$ decreases. This is in line with expectations, as the greater the action length required to complete a certain instruction, the more complex the instruction is, and the more difficult it is to complete.
@@ 5.4 Effect of fine-tuning
Table 6: Performance of fine-tuned model.
@@
Qwen-2.5-vl-7B-FT 9.7 [+6.6] 0.49 [+0.24] 12.5 [+11.0] 1.21 [+0.26] 1.17 [-0.99]
GPT-4.1 7.2 0.35 5.5 0.98 1.98
@@
To save resources, we only fine-tuned Qwen-2.5-VL-7B and adopted the parameter-efficient fine-tuning method of LoRA, with the LoRA rank set to 4. Following the method of sampling the test set, we randomly sampled 1,000 data episodes from the training set for fine-tuning. These data covered all users, and the proportion of data for each user was the same as their proportion in the training set. These 1,000 data episodes were reorganized according to the input and output formats of the two tracks, respectively. The prompts used in fine-tuning are the same as those we designed for generalist models. Finally, we trained separately on two tracks and obtained two fine-tuned models, each suitable for one of the two tracks.
The performance of the fine-tuned model on the two tracks is shown in Table 6. Despite using a smaller model and less training data, the fine-tuned model achieved significant performance improvements in all main metrics. In proactive task suggestion, compared with the best-performing generalist model GPT-4.1, our fine-tuned model achieved better performance in both $SR_{1}$ and $Sim_{1}$. In proactive task suggestion, compared with the best-performing UI-TARS-1.5-7B, our fine-tuned model had a lower success rate $SR_{2}$. We consider this acceptable because UI-TARS is a model specifically designed and extensively trained for GUI grounding and GUI control, and thus has a more general instruction execution capability. However, our fine-tuned model had a significantly higher $Sim_{2}$, indicating that the action paths it selects may not be optimal but are closer to the user’s action preferences. Overall, the model fine-tuned on our collected data demonstrated stronger proactivity and personalization capabilities, being able to utilize user-related contextual information to extract potential intent patterns and action preferences from the user’s past intents and actions, which existing models have not or find it difficult to consider.
@@ 6 Conclusions
We present FingerTip 20K, a benchmark advancing mobile GUI agents toward proactive task suggestion and personalized task execution. Our data captures longitudinal user interactions, enriched with contextual information to model user-specific patterns. Experiments reveal significant gaps in existing models’ ability to leverage such contextual information. Fine-tuning Qwen-2.5-VL-7B on our data improved suggestion success rate while better aligning actions with user preferences, demonstrating the value of user-oriented training. This work establishes critical infrastructure for developing mobile agents that anticipate user needs and adapt to user action preferences.
@@ A.1 Limitations
Our study has several limitations. Firstly, all 91 contributors live in mainland China, and mainly interact with Chinese third-party apps. The recorded linguistic habits, UI layouts and action patterns may differ markedly from other regions. Secondly, our LoRA fine-tuning uses only 1000 episodes and a single 7B model. Due to cost constraints, we did not conduct larger-scale fine-tuning experiments nor verify its generalization. Finally, we assume that screenshots can be stored and shared after anonymization. In practice, fine-grained UI traces can still contain unique visual features that allow re-identification. Techniques such as selective redaction or synthetic replay should be explored before large-scale deployment.
@@ A.3 Data collection
The data collection was carried out through crowdsourcing, and participants were paid in accordance with the living wage laws of their country. Participants were informed of the expected use of the collected data and signed a data usage agreement. They were asked not to upload any data related to private information. We provided participants with detailed guidance documents and video tutorials on how to operate the FingerTip APP for data collection. All participants went through a training phase during which they became familiar with the FingerTip APP and received targeted feedback from our manually checked data. They were encouraged to avoid using overly simplified or ambiguous language to collect clear and useful intent descriptions. It should be noted that the FingerTip APP only collects data when participants actively use it. It does not automatically collect data at other times.

### v2

@@ 4.3 Data statistics
The summary of data statistics is presented in Table 1. Additionally, Figure 4 reports the distribution of user intent length, episode length, intent categories, and app name in all data. The intent categories are determined by DeepSeek-V3 (Liu et al., 2024).
@@ 4.4 Personalized action analysis
Table 2: The action space of an agent when interacting with a mobile phone environment.
Action Parameter
@@
content="
type text="
@@
direction="
press_back -
press_home -
press_recent -
wait -
finish -
Figure 5: Personalized action analysis. We demonstrated the similarities of user action sequences in six intent categories. The similarities were higher among the same users or users of the same type, while the similarities between users of different types were lower.
To verify the personalized differences in actions among users of different types, we first simply classified users into different categories based on age groups. Then, we randomly sampled one piece of data from each of the 40 intent categories. For the action sequence of such a piece of data, we calculated the Levenshtein similarity with the action sequence of the most intent-similar data from (i) the same user, (ii) users of the same type, and (iii) users of different type. All similarities were normalized to [0, 100] and plotted in Figure 5. It can be seen that even when performing similar intents, the similarity of action sequences with users of different type is significantly lower than that of the same user or users of the same type, indicating that user preferences on action sequences do exist and are measurable.
@@ 5 Experiments
We conducted experiments on some generalist models and some GUI-control agents built on specifically designed models, evaluating their capabilities on the two tracks proposed in the FingerTip 20K benchmark and assessing their performance under different task difficulties. Additionally, we fine-tuned a model using the collected data. For details on the data splits (including the training set, validation set, and two test sets), please refer to Appendix A.5.
@@ 5.1 Experimental setup
The LLMs we experiment with in this track include GPT-4.1, Qwen-VL-Max, DeepSeek-VL2 (Wu et al., 2024) and Qwen-2.5-VL-7B (Bai et al., 2025). We also introduce Qwen-2.5-VL-72B (Bai et al., 2025) to compare with the 7B version; and Qwen-QVQ-Max (thinking model) to compare with other non-thinking models. We set the temperature to zero for all models. For proactive task suggestion, the agent only needs one query to output the predicted intent. Since this is a brand new track proposed in our benchmark, there is no mature agent design available for direct use. We have designed a simple prompt to provide to all models evaluated in this track. This prompt contains all necessary inputs (see Section 3.1 and Appendix A.7.1).
@@
In this track, in addition to the generalist models mentioned above, we also experiment with three GUI-control agents built on specifically designed models, including Aguvis-7B (Xu et al., 2024b), CogAgent-9B (Hong et al., 2024) and UI-TARS-1.5-7B (Qin et al., 2025). We also introduce AutoDroid Wen et al. (2024) and AppAgent Zhang et al. (2025), two GUI-control agents based on prompt engineering (using GPT4.1 as the base model). For personalized task execution, the agent needs to interact with the environment in multiple steps to fulfill the user’s instructions. We connect a physical phone to the computer via USB and use Android Debug Bridge (ADB) to provide this environment. Using an emulator would be a more convenient approach, but due to strict app control measures, most Chinese apps can only run on physical phones rather than emulators. For the generalist models, we designed a simple prompt to guide their output of the next action, with the action space consistent with Table 2. This prompt contains all necessary inputs (see Section 3.2 and Appendix A.7.2). For the GUI-control agents, they have specific format requirements for input and output. To ensure normal output effects, their original prompts were used, and the input information in Section 3.2 was uniformly integrated into these original prompts. Their output was converted into a form consistent with the action space.
@@
In proactive task suggestion track, the goal of the agent is to maximize the textual similarity between the output and the user’s true intent. We use a pre-trained model, paraphrase-multilingual-MiniLM-L12-v2 (Reimers and Gurevych, 2019), to convert the agent’s output and the user’s true intent into embedding vectors and calculate their cosine similarity $S_{1}$. And, we calculate the Levenshtein similarity $S_{2}$ of these two strings. Both similarities are normalized to the range of [0, 1]. Finally, we take $Sim_{1}=(S_{1}+S_{2})/2$ to comprehensively represent the text similarity. In addition to this numerical metric, we also use DeepSeek-V3 (Liu et al., 2024) to directly determine whether the agent’s output and the user’s true intent can be regarded as the same intent and provide a binary value to evaluate whether the agent successfully predicted the user’s intent, thereby calculating the success rate $SR_{1}$.
In personalized task execution track, the primary goal of the agent is to execute user instructions in a personalized manner. We calculate the final success rate $SR_{2}$ by manually checking whether the environment state when the agent outputs $finished()$ matches the user’s instructions. In addition, when the agent steps exceed 2.5 times the golden steps, the task is automatically considered a failure. Note that the path to successfully execute the user’s instructions is not unique. The agent should also make the action sequence reflect the user’s action preferences as much as possible. We do not require the agent’s action at each step to be exactly the same as the user’s golden action. Instead, we calculate the Levenshtein similarity $S_{I}$ of the agent’s complete action sequence and the user’s complete action sequence as two strings. Then, following the approach in Section 4.4, we take the data that is most similar to the current user’s intent from the users of different type, and calculate the Levenshtein similarity $S_{\text{II}}$ of the agent’s complete action sequence and this data’s complete action sequence. Finally, we take the value $Sim_{2}=S_{I}/S_{\text{II}}$. It is obvious that the larger this value is, the more similar the agent’s action sequence is to that of the current user, and the more different it is from that of users of different type. In addition, we measure execution efficiency by comparing the agent steps with the user’s golden steps to calculate the step ratio when the agent successfully execute the user’s instructions. For the two tracks, we also tallied the average time and token count consumed per query to assess the model’s cost.
@@ 5.2 Overall performance
The overall performance of the models we evaluated in proactive task suggestion is shown in Table 3. Note that here we set the number of $O$ (the first few screenshots of the user’s current actions) to 0. This makes the task quite challenging. The thinking model Qwen-QVQ-Max surpassed GPT-4.1, achieving the best performance among the generalist models with $SR_{1}=12.8$ and $Sim_{1}=0.39$, but also took the longest time and the most tokens. From $SR_{1}$, it can be intuitively seen that the success rate of all models in predicting the user’s intent is very low. Additionally, we conducted a user study where 20 human annotators (distinct from the users who collected the data) labeled a subset of the test set (a total of 400 episodes), achieving a success rate of 30.3%. This highlights the significant gap between the existing models and human in proactive task suggestion capabilities.
Table 3: Overall performance of proactive task suggestion.
@@
Qwen-2.5-VL-72B 7.0 0.31 5.45 963
Qwen-QVQ-Max 12.8 0.39 10.60 2335
Human 30.3 0.57 - -
Table 4: Overall performance of personalized task execution.
@@
The overall performance of the models we evaluated in personalized task execution is shown in Table 4. Qwen-QVQ-Max and UI-TARS-1.5-7B achieved the best performance among the generalist models and GUI-control models respectively. AppAgent achieved the best performance among all models in $Sim_{2}$ and step ratio, possibly due to its proficiency in learning from human demonstrations, but time and token costs also increased significantly. The $SR_{2}$ of the generalist models were all very low, mainly due to their lack of precise GUI grounding ability, which led to incorrect UI coordinates being output even when they could correctly analyze the next action, thus failing to interact with the environment accurately. In contrast, the GUI-control models, having undergone targeted training, had stronger abilities to execute instructions and interact with the UI environment, resulting in higher $SR_{2}$, with UI-TARS-1.5-7B having the highest at 38.5. However, the $Sim_{2}$ of all models were approximately 1, indicating that the agent’s action sequence did not favor either the current user or users of different type. This might suggest that the agent tends to complete tasks in a general way without catering to the specific action preferences of users, thus failing to complete tasks in a personalized manner.
@@ 5.3 Effect of task difficulty
Figure 6: Performance under different task difficulties. (a) The variation of $SR_{1}$ under different numbers of input screenshots. (b) The variation of $SR_{2}$ under different action lengths.
@@
For personalized task execution (see Figure 6.b), we calculated the $SR_{2}$ of GUI-control models on different subsets of action length (i.e., the number of action steps) in the test set. It can be seen that as the action length increases, the $SR_{2}$ decreases. This is in line with expectations, as the greater the action length required to complete a certain instruction, the more complex the instruction is, and the more difficult it is to complete.
@@ 5.4 Effect of fine-tuning
We fine-tuned Qwen-2.5-VL-7B and adopted the parameter-efficient fine-tuning method of LoRA, with the LoRA rank set to 4 or 64. Following the method of sampling the test set, we randomly sampled 1,000 data episodes from the training set for fine-tuning. These data covered all users, and the proportion of data for each user was the same as their proportion in the training set. We also used the complete training set (16,000 episodes) for fine-tuning. The data episodes were reorganized according to the input and output formats of the two tracks, respectively. The prompts used in fine-tuning are the same as those we designed for generalist models. Finally, we trained separately on two tracks and obtained two fine-tuned models, each suitable for one of the two tracks.
Table 5: Performance of fine-tuned models. In square brackets [X] we report the performance increase from the un-fine-tuned Qwen-2.5-VL-7B.
@@
Qwen-2.5-VL-7B-FT-1k-r4 9.7 [+6.6] 0.49 [+0.24] 12.5 [+11.0] 1.21 [+0.26] 1.17 [-0.99]
Qwen-2.5-VL-7B-FT-1k-r64 11.8 [+8.7] 0.50 [+0.25] 12.5 [+11.0] 1.26 [+0.31] 1.17 [-0.99]
Qwen-2.5-VL-7B-FT-all-r4 20.3 [+17.2] 0.52 [+0.27] 15.0 [+13.5] 1.32 [+0.37] 1.15 [-1.01]
Qwen-2.5-VL-7B-FT-all-r64 26.0 [+22.9] 0.55 [+0.30] 15.5 [+14.0] 1.42 [+0.47] 1.13 [-1.03]
Qwen-QVQ-Max 12.8 0.39 9.5 1.04 1.94
@@
The performance of fine-tuned models on the two tracks is shown in Table 5. Despite using a smaller model and less training data, the fine-tuned models achieved significant performance improvements in all main metrics. Increasing the LoRA rank or the amount of training data both improve the model’s performance, with the increase in training data having a particularly significant effect. In proactive task suggestion, compared with the best-performing generalist model Qwen-QVQ-Max, our fine-tuned models achieved better performance in both $SR_{1}$ and $Sim_{1}$. In personalized task execution, compared with the best-performing UI-TARS-1.5-7B, our fine-tuned models had a lower success rate $SR_{2}$. We consider this acceptable because UI-TARS is a model specifically designed and extensively trained for GUI grounding and GUI control, and thus has a more general instruction execution capability. However, our fine-tuned models had a significantly higher $Sim_{2}$, indicating that the action paths they select may not be optimal but are closer to the user’s action preferences. When trained on the entire training set with a LoRA rank of 64, Qwen-2.5-VL-7B outperforms all the un-fine-tuned models in the experiment in terms of $SR_{1}$, $Sim_{1}$, and $Sim_{2}$, achieving the best performance. Overall, the models fine-tuned on our collected data demonstrated stronger proactivity and personalization capabilities, being able to utilize user-related contextual information to extract potential intent patterns and action preferences from the user’s past intents and actions, which existing models have not or find it difficult to consider.
More experiments
In Appendix A.6 we conducted more experiments to study the influence of other factors.
@@ 6 Conclusions
We present FingerTip 20K, a benchmark advancing mobile LLM agents toward proactive task suggestion and personalized task execution. Our data captures longitudinal user interactions, enriched with contextual information to model user-specific patterns. Experiments reveal significant gaps in existing models’ ability to mine such patterns. Fine-tuning Qwen-2.5-VL-7B on our data improved suggestion success rate while better aligning actions with user preferences, demonstrating the value of user-oriented training. This work establishes critical infrastructure for developing mobile agents that anticipate user needs and adapt to user action preferences.
7 Ethics statement
Our data collection involves human participants. We detail our data collection process and the multiple measures we have taken to reduce the risk of privacy leakage in Appendix A.3. We also discuss the broader impacts of this study in Appendix A.2.
8 Reproducibility statement
To ensure the reproducibility of our work, we provide all the necessary resources and code used in this paper. All adopted models are fully open source or publicly accessible. Our project code, including the data format, data splits, and evaluation process of FingerTip 20K, can be publicly accessed via the following anonymous link: https://github.com/tsinghua-fib-lab/FingerTip-20K.
@@ A.1 Limitations
Our study has several limitations. Firstly, all 95 contributors live in mainland China, and mainly interact with Chinese third-party apps. The recorded linguistic habits, UI layouts and action patterns may differ markedly from other regions. Secondly, our LoRA fine-tuning uses only a single 7B model. Due to cost constraints, we did not conduct larger-scale fine-tuning experiments. Finally, we assume that screenshots can be stored and shared after anonymization. In practice, fine-grained UI traces can still contain unique visual features that allow re-identification. Techniques such as selective redaction or synthetic replay should be explored before large-scale deployment.
@@ A.3 Data collection
The data collection was carried out through crowdsourcing, and participants were paid in accordance with the living wage laws of their country. Participants consist of one-third undergraduates, one-third postgraduates, and one-third employed individuals, including 54 males and 41 females, whose ages range from 18 to 60, with an average age of 25.9. Participants filled out a questionnaire, which collected their user profiles. Participants were informed of the expected use of the collected data and signed a data usage agreement. They were asked not to upload any data related to private information. We provided participants with detailed guidance documents and video tutorials on how to operate the FingerTip APP for data collection. All participants went through a training phase during which they became familiar with the FingerTip APP. They were encouraged to avoid using overly simplified or ambiguous language to collect clear and useful intent descriptions. They were clearly informed that they should not perform redundant or useless operations during the demonstration process, and the operation speed should not be too fast to avoid frequent repetitive operations. However, minor noisy operations (e.g., users making a typo or accidentally touching advertisements) are realistic situations in human interaction. A robust agent must be able to handle such scenarios. Even if the demonstrations are not collected from daily life but by recruiting annotators to perform operations in a simulator like existing datasets, such noise cannot be completely avoided. Therefore, we allow for its existence. During data collection, we conducted multiple timed quality checks on the data submitted by each participant and manually deleted the low-quality data. We also provided quality feedback to the corresponding participants, reminding them how to submit higher-quality data.
It should be noted that the FingerTip APP only collects data when participants actively use it. It does not automatically collect data at other times. Participants can check or delete the data they upload at any time. We conducted two rounds of inspections. We first manually inspected the data and removed those that obviously involved privacy. Then, we used Qwen-VL-Max to examine the first and last screenshots of each episode and determine whether it involved privacy. Those episodes marked as potentially involving privacy were then rechecked by humans.
Our primary goal for collecting the data was to capture deep and longitudinal user interactions in daily life settings. We believe that this context-rich dataset, even from a single region, provides a crucial foundation for the novel tasks of proactive task suggestion and personalized task execution. Considering the cost, we did not collect data in other regions. To our knowledge, previous datasets such as Rawles et al. (2023); Li et al. (2024); Chen et al. (2024) also contain a single language and UI ecosystem. We believe that this is a sufficient start for a first-of-its-kind study. However, user diversity is a crucial aspect in ensuring the global generalizability of our findings. To facilitate broader research, we plan to open source our APP for data collection. It can run on any (new version) Android personal phone, providing support for data collection in other regions and languages. We believe that our data collection methods and evaluation methods are universal.